OpenStudy (anonymous):

A car is traveling at a speed of 44 feet per second. What is the car's speed in miles per hour? How many miles will the car travel in 5 hours? In your computations, use the fact that 1 mile is equal to 5280 feet. Do not round your answers.

OpenStudy (anonymous):

first let find out how many feet he is going a minute. 44*60=2640 then we'll find how many feet per hour. 2640*60=158400 then to need to convert feet to miles. 158400/5280=30

OpenStudy (anonymous):

so the car is go 30mph. in five hours it would travel 150 miles

OpenStudy (anonymous):

do you need anymore help?